OpenStudy (anonymous):

determine if the doubly infinite improper integral converges and, if so, evaluate it determine if the doubly infinite improper integral converges and, if so, evaluate it @Mathematics

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty}dx/(1+x^{2})\]

OpenStudy (bookylovingmeh):

wait this not math

OpenStudy (bookylovingmeh):

wrong grup

OpenStudy (anonymous):

yea it is...

OpenStudy (bookylovingmeh):

sorry

myininaya (myininaya):

 so we need to break up the integral i'm going to choose the constant to be 0 so we have \[\int\limits_{-\infty}^{0}\frac{1}{1+x^2} dx +\int\limits_{0}^{\infty}\frac{1}{1+x^2} dx\]

OpenStudy (anonymous):

well now i feel dumb

OpenStudy (anonymous):

The infamous arctan(x) function. We meet again.

myininaya (myininaya):

\[\lim_{a \rightarrow - \infty}[\tan^{-1}(x)]_{a}^0 +\lim_{b \rightarrow \infty}[\tan^{-1}(x)]_0^{b}\]

myininaya (myininaya):

\[\lim_{a \rightarrow -\infty}[\tan^{-1}(0)-\tan^{-1}(a)]+\lim_{b \rightarrow \infty}[\tan^{-1}(b)-\tan^{-1}(0)]\]

myininaya (myininaya):

\[0-\lim_{a \rightarrow -\infty} \tan^{-1}(a)+\lim_{b \rightarrow \infty}\tan^{-1}(b)-0\] \[0-\frac{-\pi}{2}+\frac{\pi}{2}-0\] so both integrals are convergent so that means the one we started out with his convergent and actually converges to pi

OpenStudy (anonymous):

thank you